OpenStudy (anonymous):

Solve the following system of equations. 6x+9y = -3 6x+5y = 9

OpenStudy (triciaal):

to solve by elimination subtract the second from the first solve for y then use that value in the original to find x

OpenStudy (hockeychick23):

ok so the simplest way to solve this equation is to use the process of elimination. So we subtract the second equation from the first so 6x-6x=0 and 9y-5y= 4y. -3-9=-12. So now we have 4y=-12. Do you know the next step?

OpenStudy (anonymous):

divide both sides by 4?

OpenStudy (anonymous):

x = 4 y= -3